)t takes 0.117 s for the rock to pass by a 1.80-m high window. How far from the top of the building is the top of the window? (Express your answer to three significant figures.) Submit You currently have 2 submissions for this question. Only 10 submission are allowed. You can make 8 more submissions for this question.............................................
Answer #1

h1 = height from top of building to top of window

h2 = height of window

t1 = time to travel h1

t2 = time to travel h2

if initial velocity of rock V0 = 0

h1 = (1/2)*g*t1^2

h2 = (1/2)*g*t2^2

h2-h1 = (1/2)*g*(t2^2-t1^2)

t2 -t1 = 0.117

t2 = t1+ 0.117

1.8 = (1/2)*9.8*((t1+0.117)^2-t1^2)

t1 = 1.51 s



h1 = (1/2)*9.8*1.51^2



h1 = 11.2 m